Write and equation of the line, in the form y = mx + b, that is parallel to the line y - 3x = 6, and has a y-intercept of -2.
Thank you so much for your help

Answer by Simnepi(216)

The form y = mx + b gives the equation of a straight line where m is the slope of the line(the gradient) and b is the intercept of the line and the y axis.
All lines with the same gradient (i.e. the same number in front of the x) are parallel.
First, let's rearrange your line equation into the correct form of y = mx + b since
y - 3x = 6 has the x term on the 'wrong side' of the equals sign.
Let's take it across the equals sign and change it's sign as we do so.
(it changes from a -3x into a +3x)
so now we have
y = 3x + 6
This line has a gradient of 3 and intercepts the y axis at +6
Which means that the line we are looking for must also have a gradient of 3.
The line we want has a y-axis intercept of -2.
We can now write the line's equation
y = 3x - 2